Argentina Partners is concerned about the possible effects of inflation on its operations. Presently, the company sells 77,000 units for $55 per unit. The variable production costs are $36, and fixed costs amount to $870,000. Production engineers have advised management that they expect unit labor costs to rise by 20 percent and unit materials costs to rise by 15 percent in the coming year. Of the $36 variable costs, 50 percent are from labor and 25 percent are from materials. Variable overhead costs are expected to increase by 25 percent. Sales prices cannot increase more than 10 percent. It is also expected that fixed costs will rise by 4 percent as a result of increased taxes and other miscellaneous fixed charges.

The company wishes to maintain the same level of profit in real dollar terms. It is expected that to accomplish this objective, profits must increase by 5 percent during the year.

Required:

a. Compute the volume in units and the dollar sales level necessary to maintain the present profit level, assuming that the maximum price increase is implemented.

b. Compute the volume of sales and the dollar sales level necessary to provide the 5 percent increase in profits, assuming that the maximum price increase is implemented.

c. If the volume of sales were to remain at 77,000 units, what price increase would be required to attain the 5 percent increase in profits? Calculate the new price.
Answer #1

Answer: @ Existing : Sales = $ss variable cost per unit = labor cost + material cost towthcad cost = (36 x 50%) +($.36x854.)+NW: $55 X 110.. - 60.5 variable cast per unit = labor cost + Merterial cost tovehed cost (flax 12017) + ( 9x1154)+(x last.) =. Dollar sdea =uqalacaq. os Dollar sales = $4978969, Ⓡ. Ő volumes in cristas - frange profit + fixed cost ) fie - 2 -(593,60= 62.14-55 = 9.14 inercase ofe = $7.141455 = 12.98.4